# Calibrating Verbal Uncertainty as a Linear Feature to Reduce Hallucinations

### v1

@@ Abstract
LLMs often adopt an assertive language style also when making false claims. Such “overconfident hallucinations” mislead users and erode trust. Achieving the ability to express in language the actual degree of uncertainty around a claim is therefore of great importance. We find that “verbal uncertainty” is governed by a single linear feature in the representation space of LLMs, and show that this has only moderate correlation with the actual “semantic uncertainty” of the model. We apply this insight and show that (1) the mismatch between semantic and verbal uncertainty is a better predictor of hallucinations than semantic uncertainty alone and (2) we can intervene on verbal uncertainty at inference time and reduce hallucinations on short-form answers, achieving an average relative reduction of 32%.
@@ 1 Introduction
We next leverage our findings to study hallucinations through the lens of uncertainty features. As illustrated in Figure 2, we highlight the misalignment between verbal uncertainty and the uncertainty about what meaning to convey in model outputs, i.e. semantic uncertainty, as an important factor contributing to hallucinations. We propose a novel method to detect model hallucinations that incorporates both verbal and semantic uncertainty information, and show that it outperforms detection methods that reply solely on semantic uncertainty. Next, we propose a method, Mechanistic Uncertainty Calibration (MUC), that steers LLM activations using VUF to calibrate verbal uncertainty with the model semantic uncertainty.
We demonstrate that MUC effectively reduces model hallucinations, achieving an average relative reduction of 31.9% in short-form QA tasks. It also induces nuances expressions of uncertainty and achieves a 28.0% improvement in the alignment between verbal and semantic uncertainty.
@@ 2.1 Semantic Uncertainty
Semantic Uncertainty (SU) refers to the intrinsic uncertainty of a model in the semantic meaning relative to a statement. It reflects the confidence level of a model’s prediction, focusing on semantic diversity and consistency rather than lexical diversity (Lin et al., 2022; Kadavath et al., 2022; Mielke et al., 2022).
We follow (Kuhn et al., 2023) in operationalizing this concept in terms of Semantic Entropy that can be calculated as follows: for a given question, one first samples multiple model answers and cluster them into semantically equivalent groups, within which the meaning of every answer can be inferred from each other. Semantic entropy is then computed as the entropy of the probability distribution over semantic equivalence answer classes.
@@ 2.2 Verbal Uncertainty
Verbal Uncertainty (VU) refers to the degree of uncertainty in a statement that a speaker expresses through their choice of word and sentence structures. We follow recent work in expression decisiveness quantification and employ "LLM-as-a-Judge" to measure VU (Yona et al., 2024; Zheng et al., 2023).
Specifically, we sample multiple answers for each question and prompt an auxiliary evaluator LLM to directly assign a VU score to each answer. The VU for a question is the average of VU scores of all answers.  [^1] See Appendix 7 for the prompt used for VU estimation. This approach has been shown to yield reliable uncertainty estimations that are highly correlated with human judgments of perceived assertiveness (Yona et al., 2024; Fagen-Ulmschneider, 2023). To further validate the robustness of "LLM-as-a-Judge", we compute sentence embedding cosine similarities with pre-defined prototypical uncertainty expressions and find a high correlation with VU scores returned by LLMs (see Appendix 11.2 for details).
Figure 3: Visualization of verbalized certain (blue) vs. uncertain (pink) query representations exacted from selected layers of (a) Llama-3.1-8B-Instruct, (b) Mistral-7B-Instruct-v0.3, and (c) Qwen2.5-7B-Instruct on three datasets (TriviaQA, NQ-Open, PopQA). Please refer the Appendix 11 exacted for all layers of three models.
@@ 2.3 Hallucination arise from Miscalibration between Semantic and Verbal Uncertainty
As introduced in § 1, ideally, VU should align with SU to allow the model to express uncertainty in the semantic meaning of its outputs. However, observations indicate these uncertainties are not correlated, leading to occurrences of hallucinations. In this section, we will quantitatively investigate and demonstrate the miscalibration between semantic and verbal uncertainty by analyzing samples from TriviaQA  [^2] See Appendix 10 for examples..
Following Kossen et al. (2024); Farquhar et al. (2024), for each question, we generate a response using a low temperature (0.1) to obtain the most likely answer, and then sample multiple responses using a high temperature (1.0). We categorize the samples into two primary groups based on the VU level of the most likely answer: those that include abstentions (abstained) and those that do not (complying). We further subdivide these categories. For complying responses, we assess whether the answers are hallucinated or correct (hallucinated/correct). For abstained, we determine if the model consistently refuses (i.e., all samples), or if it complies at least once among the multiple sampled answers (“partly abstained”).
As shown in Figure 2, abstained responses have high VU, which is expected. Consistently abstained ones have low SU, but this is not a problematic mismatch, rather an artifact of using semantic entropy as a proxy for SU: these are cases where the model “knows that it does not know” and deals with them appropriately. There is however a large segment of complying answers with high SU and low VU that are hallucinations  [^3] There are outlier hallucinations with low SU. For a detailed analysis, please refer to Appendix 10.3.: this is the focus of our intervention. We show in § 4.1 that combining predictions of LU and SU helps detecting hallucination. Moreover, we show in § 4.2 that modulating VU to better reflect SU is crucial to prevent confident hallucinations and optimize the trade-off balance between hallucinations and false abstention.
@@ 3.1 Feature Extraction
We first collect a set of pairs of questions and generated answers for which the model generates answers with high VU, denoted as question $x\in\mathcal{D}_{uncertain}$, and another set for questions on which the model generates certain answers with low VU, denoted as $x\in\mathcal{D}_{certain}$. We construct $\mathcal{D}_{uncertain}$ and $\mathcal{D}_{certain}$ by choosing top $N_{uncertain}$ and bottom $N_{certain}$ pairs sorted by their VU score, which is calculated using the LLM-as-a-Judge method described in § 2.2 with the Llama3.1-70B-Instruct model (Dubey et al., 2024). To generate answers, we use the “uncertainty” eliciting prompt provided in Appendix 7.1. We then calculate the difference between the model’s mean last-token residual stream activations $\mathbf{h}^{(l)}(x)$ for each layer $l$ when processing these two sets of questions, as illustrated in the following equation:
$\begin{split}\mathbf{r}_{\text{VU}}^{(l)}=&\frac{1}{|\mathcal{D}_{\text{uncertain}}|}\sum_{x\in\mathcal{D}_{\text{uncertain}}}\mathbf{h}^{(l)}(x)\\ &-\frac{1}{|\mathcal{D}_{\text{certain}}|}\sum_{x\in\mathcal{D}_{\text{certain}}}\mathbf{h}^{(l)}(x)\end{split}$ (1)
@@ 3.2 Discovery of Linear Verbal Uncertainty Features
We validate the causal connection between VUFs and the model’s VU by analyzing the generation behavior as we modulate the strength of the corresponding feature through simple linear interventions. Inspired by Li et al. (2024), we intervene on model activations of all tokens by steering them along a set of VUF directions. Observations from visualization indicate that VUFs are exacted from the middle to the last layers. For each layer $l$, we extract VUs $\mathbf{r}_{\text{VU}}^{(l)}\in\mathbb{R}^{d_{model}}$. Specifically, the VU feature vector $\mathbf{r}_{\text{VU}}^{(l)}$ serves as a directional guide for steering activations, as described in the equation below:
@@
where $\alpha$ is the intensity of intervention, and $\mathbf{r}_{\text{VU}}^{(l)}$ is the verbal uncertainty feature at layer $l$. The results presented in Figure 4 show that adding VUFs to model activations ($\alpha>0$) increases the VU of the model outputs. Conversely, removing VUFs to activations ($\alpha<0$) decreases this uncertainty. As the intensity of VUFs ($|\alpha|$) gets stronger, the VU scores exhibit greater changes. This trend remains consistent across all models and datasets we studied. This show the potential of VU calibration in model generation. We will further explore how to leverage this phenomenon in §4.2.
Interestingly, although Qwen2.5 exhibits a similar trend, it is significantly less sensitive compared to Llama3.1 and Mistral.
@@
An important question is whether the obtained VUFs are dependent on the dataset. To investigate this, we measure the cosine similarity of distinct VUFs extracted from TriviaQA, NQ-Open, and PopQA datasets respectively. We explore VUFs obtained from each layer of three models. The results presented in Figure 5 show a high cosine similarity between VUFs from different datasets, particularly in the middle and subsequent layers. This pattern is consistent across all models and datasets we examined.
In addition, as shown in Figure 6, VUFs exacted from TriviaQA can control the VU level of other datsets: NQ-Open and PopQA. Adding or removing TriviaQA VUFs increases or decreases the VU of model outputs for these datasets. These two findings indicate that VUFs are consistent across different datasets, suggesting that a universal VUF can be derived and utilized in our experiments further in the paper. Similar results using other VU scores are provided in Appendix 11.4.
Figure 6: Causal Validation on NQ-Open and PopQA with the VUF from the OOD dataset TriviaQA.
@@ 4 Verbal Uncertainty and Hallucination
Confident hallucinations arise from a miscalibration between VU and SU, where the model fails to express its high uncertainty in its generated output. By exploiting this miscalibration, we can detect hallucinations (§ 4.1). Additionally, we aim to mitigate confident hallucinations by calibrating these two uncertainties using VUFs discovered in § 3 (§ 4.2).
@@ 4.1 Hallucination Detection with Semantic and Verbal Uncertainty
AUROC ACC AUROC ACC AUROC ACC
TriviaQA 66.85 66.0 79.71 80.8 73.53 80.1
NQ-Open 54.07 53.9 66.02 70.3 57.15 71.3
PopQA 70.17 65.6 75.66 81.1 74.76 81.0
Table 2: Detection Results with Probe on Llama-3.1-8B-Instruct
Dataset
Hallucination
[…]
VU/SU
Disagree. Rate $\downarrow$
[…]
VU for
Incorrect $\uparrow$
VU for
Correct
before after before after before after before after before after before after before after
Llama
TriviaQA 23.3 19.0 71.3 70.6 5.4 8.2 21.50 21.40 0.59 0.63 0.50 0.55 0.16 0.16
NQ-Open 40.2 26.2 50.7 47.7 9.1 19.6 35.10 18.90 0.38 0.69 0.37 0.54 0.17 0.24
PopQA 33.7 21.6 23.5 21.0 42.8 55.8 50.70 44.70 0.05 0.34 0.61 0.73 0.17 0.20
Average 32.4 22.3 48.5 46.4 19.1 27.9 35.80 28.30 0.34 0.55 0.49 0.61 0.17 0.20
Mistral
TriviaQA 30.2 19.7 67.9 67.0 1.9 6.2 27.50 16.80 0.46 0.66 0.19 0.39 0.04 0.05
NQ-Open 52.2 40.8 41.7 39.4 6.1 13.7 46.80 19.80 0.24 0.58 0.23 0.40 0.07 0.10
PopQA 58.2 26.7 26.4 23.9 15.4 43.6 50.80 28.50 0.15 0.53 0.30 0.64 0.07 0.15
Average 46.9 29.1 45.3 43.4 7.8 21.2 41.70 21.70 0.28 0.60 0.20 0.50 0.06 0.10
Qwen
TriviaQA 37.9 23.4 58.6 58.1 3.5 7.5 27.10 22.20 0.57 0.59 0.43 0.51 0.14 0.14
NQ-Open 61.6 46.8 30.4 30.1 8.0 13.4 44.50 32.50 0.31 0.38 0.39 0.46 0.18 0.19
PopQA 44.8 33.6 18.1 16.4 37.1 45.3 46.80 43.00 0.08 0.22 0.69 0.75 0.21 0.21
Average 48.1 34.6 35.7 34.9 16.2 22.1 39.50 32.60 0.32 0.39 0.51 0.57 0.17 0.18
Table 3: Mitigation Results for Llama-3.1-8B-Instruct, Mistral-7B-Instruct-v0.3, and Qwen2.5-7B-Instruct. ‘Before’ represents the original generation and ‘after’ represents the generation after Mechanistic Uncertainty Calibration.
To evaluate the performance of hallucination detection, we follow (Kossen et al., 2024; Orgad et al., 2024) and adopt the area under the receiver operating characteristic curve (AUROC) as the main metric. Additionally, we use accuracy (ACC) as a reference metric.

### v2

@@ Abstract
LLMs often adopt an assertive language style also when making false claims. Such “overconfident hallucinations” mislead users and erode trust. Achieving the ability to express in language the actual degree of uncertainty around a claim is therefore of great importance. We find that “verbal [^1] In this paper, we employ the term ‘verbal’ to mean ‘pertaining to words rather than meaning or substance,’ as opposed to ‘spoken rather than written’ (refer to Merriam-Webster’s definitions: https://www.merriam-webster.com/dictionary/verbal). Readers may substitute ‘verbal uncertainty’ with ‘expressed uncertainty’ throughout the text if they find it preferable. uncertainty” is governed by a single linear feature in the representation space of LLMs, and show that this has only moderate correlation with the actual “semantic uncertainty” of the model. We apply this insight and show that (1) the mismatch between semantic and verbal uncertainty is a better predictor of hallucinations than semantic uncertainty alone and (2) we can intervene on verbal uncertainty at inference time and reduce confident hallucinations on short-form answers, achieving an average relative reduction of ~30%.  [^2] The code is available at https://github.com/facebookresearch/verbal_uncertainty_feature_calibration.
@@ 1 Introduction
We next leverage our findings to study hallucinations through the lens of uncertainty features. As illustrated in Figure 2, we highlight the misalignment between verbal uncertainty and the uncertainty about what meaning to convey in model outputs, i.e. semantic uncertainty, as an important factor contributing to hallucinations. We propose a novel method to detect model hallucinations that incorporates both verbal and semantic uncertainty information, and show that it outperforms detection methods that rely solely on semantic uncertainty. Next, we propose a method, Mechanistic Uncertainty Calibration (MUC), that steers LLM activations using VUF to calibrate verbal uncertainty with the model semantic uncertainty.
We demonstrate that MUC effectively reduces model confident hallucinations, achieving an average relative reduction of 29.6% in short-form QA tasks. It also induces nuances expressions of uncertainty and achieves a 28.4% improvement in the alignment between verbal and semantic uncertainty.
@@ 2.1 Semantic Uncertainty
Semantic Uncertainty (SU) refers to the intrinsic uncertainty of an agent in the semantic meaning expressed by a statement. It reflects the confidence level of a model’s prediction, focusing on its meaning and disregarding paraphrastic variations (Lin et al., 2022; Kadavath et al., 2022; Mielke et al., 2022).
We follow Kuhn et al. (2023) in operationalizing this concept in terms of Semantic Entropy that can be calculated as follows: for a given question, one first samples multiple model answers and clusters them into semantically equivalent groups, within which the meaning of every answer can be inferred from each other. Semantic entropy is then computed as the entropy of the probability distribution over semantic equivalence answer classes.
@@ 2.2 Verbal Uncertainty
Verbal Uncertainty (VU) quantifies the degree of doubt a speaker expresses about a proposition $P$, either explicitly or implicitly (e.g., "I doubt…", "Possibly…"). We formally define it as the complement of the subjective probability a listener would associate with $P$, conditioned on the utterance $U$ and contextual information $C$:
$\text{VU}(U\mid C)=1-Pr(P\mid U,C)$ (1)
In the specific case of short-form QA, this definition can be instantiated with $U$ being the answer given by an agent in response to a question $C$.
Answer $U_{1}$ is more verbally uncertain than $U_{2}$ if a listener would conclude that proposition $P$ is more probable based on answer $U_{2}$ than based on $U_{1}$:
$Pr(P|U_{2},C)>Pr(P|U_{1},C)$
where, e.g.:
  • $P$: "Bournemouth is the 29th largest city in England"
  • $C$: "What is the 29th largest city in England?"
  • $U_{1}$: "Hmm, maybe Bournemouth"
  • $U_{2}$: "Bournemouth"
We follow recent work in expression decisiveness quantification and employ "LLM-as-a-Judge" to measure VU (Yona et al., 2024; Zheng et al., 2023). Specifically, we sample multiple answers for each question and prompt an auxiliary evaluator LLM to directly assign a VU score to each answer. The VU for a question is the average of VU scores of all answers.  [^3] See Appendix 7 for the prompt used for VU estimation. This approach has been shown to produce reliable uncertainty estimates that are highly correlated with human judgments of perceived assertiveness (Yona et al., 2024; Fagen-Ulmschneider, 2023). To further validate the robustness of "LLM-as-a-Judge", we compute sentence embedding cosine similarities with predefined prototypical uncertainty expressions and find a high correlation with VU scores returned by LLMs (see Appendix 11.2 for details).
Figure 3: Visualization of verbalized certain (blue) vs. uncertain (pink) query representations exacted from selected layers of (a) Llama-3.1-8B-Instruct, (b) Mistral-7B-Instruct-v0.3, and (c) Qwen2.5-7B-Instruct on three datasets (TriviaQA, NQ-Open, PopQA). Please refer to Appendix 11 for the visualization of representations exacted from all layers.
@@ 2.3 Hallucination arise from Miscalibration between Semantic and Verbal Uncertainty
Ideally, VU should align with SU to let the model faithfully express uncertainty in the semantic meaning of its outputs. However, observations indicate that the two types of uncertainty are not always correlated, resulting in hallucinations. In this section, we quantitatively investigate and demonstrate the miscalibration between semantic and verbal uncertainty by analyzing samples from TriviaQA using Llama3.1  [^4] See Appendix 10 for examples..
Following Kossen et al. (2024); Farquhar et al. (2024), for each question, we generate a response using a low temperature (0.1) to obtain the most likely answer, and then sample multiple responses using a high temperature (1.0). We categorize the samples into two primary groups based on the VU level of the most likely answer: those that include abstentions (abstained) and those that do not (complying). We further subdivide these categories. For complying responses, we assess whether the answers are hallucinated or correct (hallucinated/correct). For abstained, we determine if the model consistently refuses (i.e., all samples), or if it complies at least once among the multiple sampled answers (“partly abstained”)  [^5] We obtain the MANOVA results which indicate significant differences in 4 groups with p-values ¡ 0.0001..
As shown in Figure 2, abstained responses have high VU, which is expected. Consistently abstained ones have low SU, but this is not a problematic mismatch, rather an artifact of using semantic entropy as a proxy for SU: these are cases where the model “knows that it does not know” and deals with them appropriately. There is however a large segment of complying answers with high SU and low VU that are hallucinations  [^6] There are outlier hallucinations with low SU. For a detailed analysis, please refer to Appendix 10.3.: this is the focus of our intervention. We show in § 4.1 that combining predictions of VU and SU helps detecting hallucination. Moreover, we show in § 4.2 that modulating VU to better reflect SU is crucial to prevent confident hallucinations and optimize the trade-off balance between hallucinations and false abstention.
@@ 3.1 Feature Extraction
We first collect a set of pairs of questions and generated answers for which the model generates answers with high VU, denoted as question $x\in\mathcal{D}_{uncertain}$, and another set for questions on which the model generates certain answers with low VU, denoted as $x\in\mathcal{D}_{certain}$. We construct $\mathcal{D}_{uncertain}$ and $\mathcal{D}_{certain}$ by choosing top $N_{uncertain}$ and bottom $N_{certain}$ pairs sorted by their VU score, which is calculated using the LLM-as-a-Judge method described in § 2.2 with the Llama3.1-70B-Instruct model (Dubey et al., 2024). To generate answers, we use the “uncertainty” eliciting prompt provided in Appendix 7.1. We then calculate the difference between the model’s mean last-token residual stream activations $\mathbf{h}^{(l)}(x)$ for each layer $l$ when processing these two sets of questions. This difference is then L2 normalized, as illustrated in the following equation:
$\begin{split}\hat{\mathbf{r}}_{\text{VU}}^{(l)}=&\,\frac{1}{|\mathcal{D}_{\text{uncertain}}|}\sum_{x\in\mathcal{D}_{\text{uncertain}}}\mathbf{h}^{(l)}(x)\\ &-\frac{1}{|\mathcal{D}_{\text{certain}}|}\sum_{x\in\mathcal{D}_{\text{certain}}}\mathbf{h}^{(l)}(x)\end{split}$ (2)
$\mathbf{r}_{\text{VU}}^{(l)}=\frac{\hat{\mathbf{r}}_{\text{VU}}^{(l)}}{\|\hat{\mathbf{r}}_{\text{VU}}^{(l)}\|}$ (3)
@@ 3.2 Discovery of Linear Verbal Uncertainty Features
Effective Layer Selection
To identify the effective layers of VUFs, we analyze VUFs obtained from each layer of three different models. We measure the cosine similarity of distinct VUFs extracted from TriviaQA, NQ-Open, and PopQA datasets respectively. The results presented in Figure 4 show a high cosine similarity between VUFs from different datasets, particularly in the middle and subsequent layers. This pattern is aligned with visualization and consistent across all models and datasets we examined. Observations from both visualization and similarity across datasets indicate that reliable VUFs are best extracted from the middle to the last layers.
@@
We validate the causal connection between VUFs and the model’s VU by analyzing the generation behavior as we modulate the strength of the corresponding feature through simple linear interventions. Inspired by Li et al. (2024), we intervene on model activations of all tokens by steering them along a set of VUF directions.
For each layer $l$, we extract VUs $\mathbf{r}_{\text{VU}}^{(l)}\in\mathbb{R}^{d_{model}}$. Specifically, the VU feature vector $\mathbf{r}_{\text{VU}}^{(l)}$ serves as a directional guide for steering activations, as described in the equation below:
@@
where $\alpha$ is the intensity of intervention, and $\mathbf{r}_{\text{VU}}^{(l)}$ is the verbal uncertainty feature at layer $l$. The results presented in Figure 5 show that adding VUFs to model activations ($\alpha>0$) increases the VU of the model outputs. Conversely, removing VUFs from activations ($\alpha<0$) decreases this uncertainty. As the intensity of VUFs ($|\alpha|$) gets stronger, the VU scores exhibit greater changes. This trend remains consistent across all models and datasets we studied. This show the potential of VU calibration in model generation. We will further explore how to leverage this phenomenon in §4.2.
Interestingly, although Qwen2.5 exhibits a similar trend, it is significantly less sensitive compared to Llama3.1 and Mistral. This is due to the normalization of the VUF. Qwen embeddings have larger norms, resulting in longer distances between clusters.
To address potential circularity concerns due to the double use of LLM in VUF extraction and VU evaluation, we validate our findings using an alternative VUF extraction method, detailed in Appendix 11.5.
@@
To investigate the generalization of VUF across datasets, we use VUFs extracted from TriviaQA to control the VU level of other datasets: NQ-Open and PopQA. As shown in Figure 6, adding or removing TriviaQA VUFs increases or decreases the VU of model outputs for these datasets. These two findings indicate that VUFs are consistent across different datasets, suggesting that a universal VUF can be derived and utilized in our experiments further in the paper. Similar results using other VU scores are provided in Appendix 11.4.
Therefore, once we identify the appropriate layers for each model, these selections remain consistent across different datasets, eliminating the need to repeat the selection process for other datasets.
Figure 6: Causal Validation on NQ-Open and PopQA with the VUF exacted from the OOD dataset TriviaQA.
@@ 4 Verbal Uncertainty and Hallucination
Confident hallucinations arise from a miscalibration between VU and SU, where the model fails to express its high uncertainty in its generated output. Taking advantage of this miscalibration, we can detect hallucinations (§ 4.1). Furthermore, we mitigate confident hallucinations by calibrating these two uncertainties using VUFs discovered in § 3 (§ 4.2).
@@ 4.1 Hallucination Detection with Semantic and Verbal Uncertainty
AUROC ACC AUROC ACC AUROC ACC AUROC ACC
TriviaQA 66.85 66.0 64.83 53.5 79.71 80.8 73.53 80.1
NQ-Open 54.07 53.9 56.29 49.3 66.02 70.3 57.15 71.3
PopQA 70.17 65.6 59.33 44.9 75.66 81.1 74.76 81.0
Table 2: Detection Results on Llama-3.1-8B-Instruct. ‘Calculated’ means that the SE feature is computed after sampling multiple answers, ’Probe-Predicted’ means that SE is as predicted by a probe that takes as input the embeddings of the last token of the question, and therefore does not require sampling.
Dataset Hallucination Rate↓ Correct. Rate↑ Refusal Rate VU/SU Disagree. Rate ↓ Correlation↑ VU for Incorrect ↑ VU for Correct
Conf./Overall Conf. Overall
before after after before after before after before after before after before after before after
Llama3.1-8B
TriviaQA 23.3 19.0 21.2 71.3 70.6 5.4 8.2 21.50 21.40 0.59 0.63 0.50 0.55 0.16 0.16
NQ-Open 40.2 26.2 32.7 50.7 47.7 9.1 19.6 35.10 18.90 0.38 0.69 0.37 0.54 0.17 0.24
PopQA 33.7 21.6 23.2 23.5 21.0 42.8 55.8 50.70 44.70 0.05 0.34 0.61 0.73 0.17 0.20
Average 32.4 22.3 25.7 48.5 46.4 19.1 27.9 35.80 28.30 0.34 0.55 0.49 0.61 0.17 0.20
Mistral-7B
TriviaQA 30.2 19.7 26.8 67.9 67.0 1.9 6.2 27.50 16.80 0.46 0.66 0.19 0.39 0.04 0.05
NQ-Open 52.2 40.8 46.9 41.7 39.4 6.1 13.7 46.80 19.80 0.24 0.58 0.23 0.40 0.07 0.10
PopQA 58.2 26.7 32.5 26.4 23.9 15.4 43.6 50.80 28.50 0.15 0.53 0.30 0.64 0.07 0.15
Average 46.9 29.1 35.4 45.3 43.4 7.8 21.2 41.70 21.70 0.28 0.60 0.20 0.50 0.06 0.10
Qwen2.5-7B
TriviaQA 37.9 23.4 34.4 58.6 58.1 3.5 7.5 27.10 22.20 0.57 0.59 0.43 0.51 0.14 0.14
NQ-Open 61.6 46.8 56.5 30.4 30.1 8.0 13.4 44.50 32.50 0.31 0.38 0.39 0.46 0.18 0.19
PopQA 44.8 33.6 38.3 18.1 16.4 37.1 45.3 46.80 43.00 0.08 0.22 0.69 0.75 0.21 0.21
Average 48.1 34.6 43.1 35.7 34.9 16.2 22.1 39.50 32.60 0.32 0.39 0.51 0.57 0.17 0.18
Llama3.1-70B
TriviaQA 12.1 10.1 11.8 87.0 86.8 0.9 1.4 7.5 7.1 0.71 0.80 0.29 0.35 0.06 0.07
NQ-Open 35.7 32.3 34.0 60.8 59.5 3.5 6.5 21.1 15.1 0.49 0.73 0.27 0.36 0.08 0.09
PopQA 41.4 28.0 35.2 44.6 42.4 14.0 22.4 22.2 14.8 0.59 0.75 0.48 0.62 0.17 0.18
Average 29.7 23.5 27.0 64.1 62.9 6.1 10.1 16.9 12.3 0.60 0.76 0.35 0.44 0.10 0.11
Table 3: Mitigation Results for Llama-3.1-8B-Instruct, Mistral-7B-Instruct-v0.3, Qwen2.5-7B-Instruct, and Llama-3.1-70B-Instruct. ‘Before’ represents the original generation and ‘after’ represents the generation after Mechanistic Uncertainty Calibration. The original generation is always confident, so there is no difference between ‘Confident’ and ‘Overall’.
To evaluate the performance of hallucination detection, we follow Kossen et al. (2024); Orgad et al. (2024) and adopt the area under the receiver operating characteristic curve (AUROC) as the main metric. Additionally, we use accuracy (ACC) as a reference metric.